# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ 9.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (9.1) and (9.2), respectively [12]. In these expressions, $k$ denotes the time index while $n$ represents the frequency index. The frequency index and time index are not necessarily associated with physical time unit or frequency unit, respectively; rather, the naming convention provides a convenient way to distinguish between the two opposite domains in DFT: the waveform domain $h[k]$ and the frequency domain $H[n]$.
@@
9.2 Modeling Errors of DFT
@@
Although mathematically self-consistent, DFT implies a periodic boundary condition in the time domain [29]. Consequently, it is only an approximation for the mathematically defined continuous Fourier transform (FT) [29]. For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [12]. Figure 34 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency. If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise. Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [12].
@@ 9.3 FFT (Fast Fourier Transformation)
When $N=2^{L}$, where $L$ is a natural number, the generalized Danielson-Lanczos lemma [12] can be applied to DFT to produce FFT [12].
  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations as Formula (2.14), (2.15), (2.16), and (2.17).
@@
  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of multiplications.
  • Each step in the forward transformation doubles the variance, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$. Consequently, the uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
@@
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zeros, causing its uncertainties to grow more rapidly.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zeros) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
@@ 9.4 Testing Signals
Empirically, except when using trigonometric library directly for FFT transformations of clean Sin and Cos signals:
  • The results obtained from Sin and Cos signals are statistically indistinguishable.
  • Similarly, the results from Sin and Cos signals at different frequencies does not show significant differences statistically except in the situation of numerical error resonance.
Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
@@ 9.5 Trigonometric Library Errors
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
@@
The constructed indexed $\sin(x)$ is referred to as the Quart indexed sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
Because the the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides numerical accuracy compared to Library function.
  • Figure 35 shows that the value difference between the Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
  • Figure 36 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but are $10^{2}$ times larger than those observed for $\sin(x)$. Therefore, the linear spectrum in Formula (9.3) may contain significant numerical errors when computed using library sine functions.
9.6 Refine Quart Sine Function
Figure 37: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
Figure 38: The value error histogram for the Quart sine function when it is calculated using 128-bit floating-point arithmetic and cast to 64-bit floating-point representation.
If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the resulting indexed sine function is referred to as the Prec sine function.
  • Figure 37 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting. As suggested in Figure 37, the Prec sine function has about $5\%$ more non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors, and $14\%$ non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors exceeding the corresponding uncertainties. These difference indicate that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic.
  • Figure 38 demonstrates that the value errors of the Prec sine function after normalized by the corresponding value ULP is uniform distributed with in the half bit of ULP. Because variance arithmetic assumes that the value error is uniformly distributed within ULP of each value, the Quart sine function overestimates value errors by one times.
@@ 9.7 Using Quart Sine for Sin/Cos Signals
Figure 39: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 40: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 41: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function. The variance arithmetic is implemented using python.
@@
Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 39 displays the value error spectrum for the forward transformation, while Figure 40 presents the value error waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.60$. In the reverse transformation, the error deviation is $0.55$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
  • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation. Even with this difference, the uncertainty tracks the corresponding value error effectively.
  • The value error waveform for the roundtrip transformation is qualitatively similar to that in Figure 40 with a similar error deviation of $0.55$.
Figure 41 illustrates that the forward transformation differs from the reverse transformation significantly.
  • In the forward transformation, error deviations reach their stable values around $0.65$ rapidly once the FFT order $L\geq 6$.
  • In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching their stable values around $0.90$ only when $L\geq 12$. This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
  • At FFT order $18$, Figure 42 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, confirming the larger error deviation.
The value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation. Despite these differences, error deviations are $0.65$ and $0.90$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
9.8 Sensitivity to Quart Sine Functions
Figure 43: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function. The variance arithmetic is implemented using C++.
Figure 44: Histograms of normalized errors between the Quart sine functions of C++, Java, and Python in the unit of least significant bits.
The reverse transformation of Sin/Cos signals is very sensitive to the underlying Quart sine function. If variance arithmetic is implemented using C++ instead of Python, Figure 43 shows that the error deviations of the reverse transform quickly reaches the stable values around $0.89$, while the forward transformation remains similar to its counterpart in Figure 41. Figure 43 also shows identical uncertainty means to their counterparts in Figure 41. Figure 44 shows that the difference of the Quart sine functions between Python and C++ is only about $1\%$ LSBs. Figure 44 also reveals that the difference of the Quart sine functions between Java and C++ is only about $0.1\%$ LSBs. Indeed, the results between Java and C++ are much more similar. Because the Python implementation gives smaller error deviations, its Quart sine function has less numerical errors on average. For discussion simplicity, the rest results are only based on the python implementation.
@@ 9.9 Using Library Sine for Sin/Cos Signals
Figure 45: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 46: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 47: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
@@
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 45 displays the value error spectrum for the forward transformation, while Figure 46 presents the vale error waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 41 and 47.
  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 35.
  • When compared with Figure 42, Figure 48 also confirms that at FFT order $18$, the normalized errors distribution is significantly broader when using the Quart sine function.
  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 47 that these in Figure 41.
  • The stable error deviations using the Library sine function are much larger than their counterparts using Quart library functions: $10.1>0.65$ for the forward transformation, $6.3>0.90$ for the reverse transformation, and $1.47>0.94$ for the roundtrip transformation.
Using the Library sine function, variance arithmetic barely achieves proper coverage for sine and cosine signals.
9.10 Numerical Error Resonance Using Library Sine
Figure 49: FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 50: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 51: Error deviation (z-axis) of FFT forward transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 51. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure 49 shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50. Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40. Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency. Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.11 Using Quart Sine for Linear Signal
Figure 53: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Quart sine function.
@@
Figure 53 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 41. Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized errors in Figure 54 versus the corresponding histograms in Figure 42. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@ 9.12 Using Library Sine for Linear Signal
Figure 55: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
@@
As shown by the difference between Figure 35 and 36, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional unspecified errors.
Figure 55 shows that proper coverage can no longer be achieved, as the value error increases faster than the uncertainty with rising FFT order. Figure 56 demonstrates that the normalized error distribution for the reverse transformation is no longer effectively bounded within three standard deviations. Although the normalized error distribution for the forward transformation using the Library sine function appears visually similar to that obtained using the Quart sine function in Figure 56, a detailed analysis reveals that the normalized error distribution contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal of the normalized errors, respectively. The presence of these extreme values is consistent with those delta-like numerical errors observed in Figure 36.
The observed difficulty of variance arithmetic in tracking Linear signals computed using the Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified numerical errors. A potential solution is to develop a new sine function library implemented directly within the variance arithmetic framework, ensuring that all numerical calculations are performed with explicit uncertainty tracking.
9.13 Ideal Coverage
Figure 57: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Figure 58: Histograms of normalized errors for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
@@ 9.13 Ideal Coverage
Code Signal Sine function Forward Reverse Roundtrip
Python Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
Java, C++ Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
Python Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
Java, C++ Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-13}$
Python Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
Java, C++ Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
Python Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
Java, C++ Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$

### v17

@@ 10.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (10.1) and (10.2), respectively [5]. As a convention, $k$ denotes time index for $h[k]$ as a waveform, while $n$ represents frequency index for $H[n]$ as a spectrum.
@@
For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [5]. Figure 23 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency. If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise. Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics
@@ 10.2 FFT (Fast Fourier Transformation)
When $N=2^{L}$, where $L$ is a natural number named FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formula (2.10), (2.11), (2.12), and (2.13).
@@
  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
  • Each step in the forward transformation doubles the variance, so the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty deviation decreases with $L$ as $\sqrt{1/2}^{L}$. Consequently, the uncertainty deviation for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
@@
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly by cancellation.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly by spreading.
The question is whether variance arithmetic can work properly in these two contrary cases.
@@ 10.4 Trigonometric Library Errors
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
@@
The constructed indexed $\sin(x)$ is referred to as the Quart sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
Because the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides superior numerical accuracy compared to the Library sine function.
  • Figure 24 shows that the absolute value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
  • Figure 25 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$ but are $10^{2}$ times larger than those observed for $\sin(x)$. Therefore, the linear spectrum in Formula (10.3) contains significantly larger numerical errors when computed using library sine functions.
@@ 10.5 Using Quart Sine for Sin/Cos Signals
With FFT order as the specific dimension, the error deviations for forward and reverse transformations resemble Figure 14, while those for round-trip transformation is almost identical to Figure 15, independent of Sin or Cos signals, or frequency of the signals. Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals. When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ due to insufficient sample count of $2^{L}$. Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
When $L=18$ and $\delta x=0$, Figure 26 shows that the error distributions of Sin/Cos signal resemble Normal distribution, with an additional Delta-like distribution at $\tilde{z}=0$. The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in sine function.
@@ 10.6 Using Library Sine for Sin/Cos Signals
Figure 28: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between uncertainty and value error. The x-axis represents the time index, while the y-axis represents both uncertainty and value error.
Figure 30: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
@@
With FFT order as the specific dimension, the error deviations for the forward and the reverse transformations resemble Figure 14, while those for the round-trip transformation is almost identical to Figure 15. In addition, the error deviations are larger than $1$ when $\delta x<10^{-15}$ for the forward transformation and $\delta x<10^{-14}$ for the reverse transformation.
When $\delta x=0$, the error deviations for the reverse transformation increase with FFT order, to $6.2$ at FFT order $18$. As shown in Figure 24, the Library sine function contains more numerical errors, so the error distribution for the reverse transformation using Library sine function in Figure 27 is more structured and broader than that in Figure 26 using Quart sine function, while the error distribution for the forward transformations is more similar. This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, but a comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 28, 29, and 30 presents the value errors for the forward, the reverse, and the round-trip transformations, respectively. In the reverse transformation, value errors exhibit a clear trend of increasing with the time index. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 30, as well as when using the Quart sine function. Figure 31 demonstrates that the error deviations increase with sine or cosine frequency. Figure 24 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern in Figure 29. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress these numerical error resonances, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals. Such resonance of numerical errors can easily and mistakenly be taken as signals.
@@ 10.7 Using Quart Sine for Linear Signal
Figure 32 and 33 show the error deviation for the forward and the reverse transformations, respectively. The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation. In other areas, both transformations achieve proper coverage with error values around $1$.
When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 34 is narrower than that in Figure 27. The corresponding error deviations are $1.5<6.2$, respectively.
@@ 10.8 Using Library Sine for Linear Signal
Figure 35 shows that the reverse error distribution when $\delta x=0$ is no longer bound. The difference between Figure 35 and Figure 34 is consistent with the large numerical errors as demonstrated in Figure 25. Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine.
Figure 36 and 37 shows a much smaller ideal coverage areas than those in Figure 32 and 33. Because result deviations grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Outside the ideal coverage region, proper coverage cannot be achieved for the reverse transformation. Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
Figure 38 shows that, even when variance arithmetic can no longer effectively track the value errors for the forward and the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as the plateau region at high $L$ and low $\delta x$ in Figure 38. Such error cancellation is due to dependence tracing in statistical Taylor expansion.
10.9 Ideal Coverage
Figure 39: Result error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Adding noise to the input can suppress unspecified input errors. After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, while the error distribution for the round-trip transformation becomes Delta. Figure 39 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
  • As expected, the resulting uncertainty deviations for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
  • As expected, the resulting uncertainty deviations for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
  • As expected, the resulting uncertainty deviations for the round-trip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
  • As expected, the resulting error deviations for the forward and the reverse transformations remain constant at $1$.
  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase with FFT order $L$ exponentially due to increasing calculation.
Signal Sine function Forward Reverse Roundtrip
Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$